# Achieving Linear Speedup with ProxSkip in Distributed Stochastic Optimization

### v3

@@ XII Proof of Lemma 5
Following a similar argument for $Q_{2,i}^{-1}$, and using $p^{2}(1+p)^{2}(\nu_{i}-c_{i})=p^{2}(1+p)^{2}(1-\frac{1}{2\chi}(1-\lambda_{i}))-(1-p)^{2}\geq 4p^{3}-\frac{4p^{2}(1-\lambda_{n})}{2\chi}\geq\frac{2p^{2}(1+\lambda_{n})}{\chi}$, we have
$\|(Q_{2,i}^{\mathrm{s}})^{-1}\|^{2}\leq\frac{2\chi}{p^{2}(1+p)^{2}(\nu_{i}-c_{i})}\leq\frac{\chi^{2}}{p^{2}(1+\lambda_{n})}.$
@@
$\|{\bf{\Gamma}}\|=\sqrt{1-\frac{1}{2\chi}(1-\lambda_{2})}<1.$
@@
XIII Proof of Lemma 6
@@
Since ${\bf{E}}^{t}=\frac{(\theta_{t}-1)}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, $\mathop{\rm Prob}(\theta_{t}=1)=p$, and $\mathop{\rm Prob}(\theta_{t}=0)=1-p$, we have
@@
$=\frac{1-p}{4\chi^{2}}\Big{(}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}+\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\Big{)}$
$=\frac{1-p}{4\chi^{2}}\Big{(}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}(\hat{{\bf{Z}}}^{t}-{\bf{X}}^{\star})\|_{\mathrm{F}}^{2}+\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}(\hat{{\bf{Z}}}^{t}-{\bf{X}}^{\star})\|_{\mathrm{F}}^{2}\Big{)}$
$\leq\frac{(1-p)(2+p^{2})}{2\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]\leq\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{(1-p)(2+p^{2})}{2\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{(1-p)(2+p^{2})}{2\chi^{2}}\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right].$ (81)
@@
$\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=\mathbb{E}\!\left[\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$=\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}+\upsilon^{2}\alpha^{2}\mathbb{E}\!\left[\left\|({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}+\frac{(p^{2}+2\chi^{2})n\alpha^{2}\sigma^{2}}{2\chi^{2}}\ .$
The last inequality holds due to $\|\hat{{\bf{\Lambda}}}_{a}\|\leq 1$, $\|\hat{{\bf{\Lambda}}}_{b}\|^{2}\leq 2$, and $\upsilon=\|({\bf{Q}}^{\mathrm{s}})^{-1}\|$. For any vectors ${\bf{a}}$ and ${\bf{b}}$, it holds from Jensen’s inequality that $\|{\bf{a+b}}\|^{2}\leq\frac{1}{\theta}\|{\bf{a}}\|^{2}+\frac{1}{1-\theta}\|{\bf{b}}\|^{2}$ for any $\theta\in(0,1)$. Therefore, letting $\theta=\|{\bf{\Gamma}}\|:=\gamma$, it holds that
$\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}$
@@
Therefore, substituting (82) and (XIII) into (81), we can conclude (47). $\Box$
XIV Proofs of Theorem 6
@@
Then, it follows from (XIV) and (87) that
@@
Since $0<\alpha<\frac{2}{L}$, $10<\frac{1-\lambda_{2}}{2}<$ 1 and $0<p^{2}\leq 1$, we have $10<\zeta<$ 1. It follows from $\Psi^{t}=\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}+\frac{2\chi\alpha^{2}}{p^{2}}\|{\bf{U}}^{t}-{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}$ that
@@
$1-\frac{p^{2}}{2\chi\kappa_{w}}=1-\frac{p^{2}(1-\lambda_{2})}{2\chi}\leq\sqrt{1-\frac{p^{2}(1-\lambda_{2})}{2\chi}}<1,\text{ and }\gamma=\sqrt{1-\frac{1-\lambda_{2}}{2\chi}}.$
Since $\tilde{\gamma}_{\mathrm{s}}=\gamma+\frac{3(1-p)(2+p^{2})}{\chi^{2}}=\sqrt{1-\frac{1-\lambda_{2}}{2\chi}}+\frac{3(1-p)(2+p^{2})}{\chi^{2}}$, we have
$\chi\geq\frac{36}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}_{\mathrm{s}}\leq\sqrt{1-\frac{p^{2}}{2\chi\kappa_{w}}}=\sqrt{1-\frac{p^{2}(1-\lambda_{2})}{2\chi}}<1.$
@@
Combining (XIV) and (91), and using $11-\tilde{\gamma}_{\mathrm{s}}^{t+1}<$ 1, it gives that

### v4

@@ XVII Proof of Lemma 5
Following a similar argument for $Q_{2,i}^{-1}$, and using $p^{2}(1+p)^{2}(\nu_{i}-c_{i})=p^{2}(1+p)^{2}(1-\frac{p}{2\chi}(1-\lambda_{i}))-(1-p)^{2}\geq 4p^{3}-\frac{4p^{2}(1-\lambda_{n})}{2\chi}\geq\frac{2p^{2}(1+\lambda_{n})}{\chi}$, we have
$\|(Q_{2,i}^{\mathrm{s}})^{-1}\|^{2}\leq\frac{2\chi}{p^{2}(1+p)^{2}(\nu_{i}-c_{i})}\leq\frac{\chi^{2}}{p(2\chi+\lambda_{n}-1)}\leq\frac{\chi^{2}}{p^{2}(1+\lambda_{n})}.$
@@
$\|{\bf{\Gamma}}\|=\sqrt{1-\frac{p}{2\chi}(1-\lambda_{2})}<1.$
@@
XVIII Proof of Lemma 6
@@
Since ${\bf{E}}^{t}=\frac{(\theta_{t}-p)}{2\chi}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}$, $\mathop{\rm Prob}(\theta_{t}=1)=p$, and $\mathop{\rm Prob}(\theta_{t}=0)=1-p$, we have
@@
$=\frac{p(1-p)^{2}}{4\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}+\frac{(1-p)p^{2}}{4\chi^{2}}\|\upsilon p({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}{\bf{W}}_{b}\hat{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}$
$\leq\frac{p(1-p)(2+p^{2})}{2\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{F}^{t}\right]\leq\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}+\frac{p(1-p)(2+p^{2})}{2\chi^{2}}\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}.$
@@
$\mathbb{E}\!\left[\|\mathcal{E}_{\mathrm{s}}^{t+1}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$ $\leq\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]+\frac{p(1-p)(2+p^{2})}{2\chi^{2}}\mathbb{E}\!\left[\|\upsilon({\bf{Q}}^{\mathrm{s}})^{-1}\hat{{\bf{P}}}^{\sf T}\widetilde{{\bf{Z}}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right].$ (82)
@@
$\mathbb{E}\!\left[\|\mathbb{G}_{\mathrm{s}}^{t}\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]=\mathbb{E}\!\left[\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star})+{\bf{S}}^{t})\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$=\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}+\upsilon^{2}\alpha^{2}\mathbb{E}\!\left[\left\|({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}{\bf{S}}^{t}\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}\;|\;\mathcal{G}^{t}\right]$
$\leq\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}+\frac{(p^{2}+2\chi^{2})n\alpha^{2}\sigma^{2}}{2\chi^{2}}\ .$
The last inequality holds due to $\|\hat{{\bf{\Lambda}}}_{a}\|\leq 1$, $\|\hat{{\bf{\Lambda}}}_{b}\|^{2}\leq 2$, and $\upsilon=1/\|({\bf{Q}}^{\mathrm{s}})^{-1}\|$. For any vectors $\bf{a}$} and $\bf{b}$}, it holds from Jensen’s inequality that $\|{\bf{a+b}}\|^{2}\leq\frac{1}{\theta}\|{\bf{a}}\|^{2}+\frac{1}{1-\theta}\|{\bf{b}}\|^{2}$ for any $\theta\in(0,1)$. Therefore, letting $\theta=\|{\bf{\Gamma}}\|:=\gamma$, it holds that
$\left\|{\bf{\Gamma}}\mathcal{E}_{\mathrm{s}}^{t}-\upsilon\alpha({\bf{Q}}^{\mathrm{s}})^{-1}\left[\begin{array}[]{c}\hat{{\bf{\Lambda}}}_{a}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \frac{p^{2}}{2\chi}\hat{{\bf{\Lambda}}}_{b}\hat{{\bf{P}}}^{\sf T}(\nabla F({\bf{X}}^{t})-\nabla F({\bf{X}}^{\star}))\\ \end{array}\right]\right\|_{\mathrm{F}}^{2}$
@@
Therefore, substituting (83) and (XVIII) into (82), we can conclude (48). $\Box$
XIX Proofs of Theorem 6
@@
Then, it follows from (XIX) and (88) that
@@
Since $0<\alpha<\frac{2}{L}$, $0<\frac{1-\lambda_{2}}{2}<1$ and $0<p^{2}\leq 1$, we have $0<\zeta<1$. It follows from $\Psi^{t}=\|{\bf{X}}^{t}-{\bf{X}}^{\star}\|_{\mathrm{F}}^{2}+\frac{2\chi\alpha^{2}}{p^{2}}\|{\bf{U}}^{t}-{\bf{U}}^{\star}\|_{\mathrm{F}}^{2}$ that
@@
[…]
Since $\tilde{\gamma}_{\mathrm{s}}=\gamma+\frac{3p(1-p)(2+p^{2})}{\chi^{2}}=\sqrt{1-\frac{p(1-\lambda_{2})}{2\chi}}+\frac{3p(1-p)(2+p^{2})}{\chi^{2}}$, we have
$\chi\geq\frac{36}{1-\lambda_{2}}\Longrightarrow\tilde{\gamma}_{\mathrm{s}}\leq\sqrt{1-\frac{p^{2}(1-\lambda_{2})}{2\chi}}<1.$
@@
Combining (XIX) and (92), and using $1-\tilde{\gamma}_{\mathrm{s}}^{t+1}<1$, it gives that